Two bulbs B1(40W,220V) and B2(100W,220V) are connected in series to an emf of 220 V. Which bulb will glow brighter?
1
Expert's answer
2012-07-05T10:11:39-0400
Let's find out the resistances of each bulb from the data given:
R1 = U1^2 /
P1 = 220^2 / 40 = 1210 Ohm
R2 = U2^2 / P2 = 220^2 / 100 = 484 Ohm
Let's
use P = I^2 * R for each bulb and find out if it is greater than
its power
rating. The one with higher power rating will get fused.
P1 = (220 / (1210 +
484))^2 * 1210 = 20.408 W
P2 = (220 / (1210 + 484))^2 * 484 = 8.163
W.
Answer: first bulb will glow brighter.